OpenStudy (anonymous):

The speed of a passenger train is 16mph faster than the speed of the freight train. The passenger train travels 310 miles in the same time it takes the freight train to travel 230 miles. Find the speed of each train.

OpenStudy (anonymous):

\[\frac{310}{x+16}=\frac{230}{x}\]310x=230(x+16) 310x= 230x +3680 310x-230x=3680 80x = 3680 x =3680/80 x =46 passenger train is : 46+16 =62 freight train : 46